OpenStudy (anonymous):

Find the sum of the geometric sequence. 1, 1/2, 1/4, 1/8, 1/16 31/16 1/24 93 1/6

OpenStudy (anonymous):

1/16 + 2/16 + 4/16 + 8/16 + 16/16

OpenStudy (tkhunny):

\(\dfrac{1 - \dfrac{1}{32}}{1 - \dfrac{1}{2}}\) You don't HAVE to do it the hard way.

OpenStudy (anonymous):

so its 31/16???

OpenStudy (tkhunny):

Any way you do it. What method did you select?

OpenStudy (anonymous):

the first one but both were good

OpenStudy (tkhunny):

Unique Answers don't care how you find them! Good work.